could you please explain how too solve this kind of equation? If you can please do: 4+3(i+ 1)=4i + 6 thankyou so much.... MLM
Answer by sdmmadam@yahoo.com(530)

Equations/27523 (2006-02-15 17:41:37): could you please explain how too solve this kind of equation? If you can please do: 4+3(i+ 1)=4i + 6 thankyou so much.... MLM
4+3(i+ 1)=4i + 6
4 +(3i+3) =4i +6
4 + (3 +3i) = 4i +6 (using additive commutativity in the field C of complex numbers)
(4+3) +3i = 4i +6 (by additive associativity in the field C of complex numbers)
7 +3i = 4i +6
7-6 = 4i - 3i (transposing, that is change side, then change sign)
1 = i
That is i = 1
which answer is RIGHT ONLY if i is just any variable x or y or a and NOT the conventional (sqroot of (-1)) in which case it cannot be equal to 1